# Soliton solutions to the Sawada–Kotera equation

### v1

@@ 5 The inverse scattering problem in the reflectionless case
In this section we present the basic results for the inverse problem for (2.8) in the reflectionless case. In that case, the inverse problem consists of the determination of the potentials $Q$ and $P$ when the transmission coefficients $T_{\text{\rm{l}}}(k)$ and $T_{\text{\rm{r}}}(k)$ are known. In general, the potentials $Q$ and $P$ cannot be determined uniquely when we only know the transmission coefficients. For the unique recovery of the potentials, we need some additional information. The specification of a dependency constant or a normalization constant at each bound state is a conventional way to recover the potentials uniquely. A bound-state dependency constant and a bound-state normalization constant are related [3] to each other via the residue of a transmission coefficient at the bound-state pole. The use of bound-state dependency constants is convenient when the potentials are recovered from the solution to a related Riemann-Hilbert problem. The use of bound-state normalization constants is appropriate when the potentials are recovered from the solution to a related linear integral equation, such as a Marchenko integral equation [7, 8, 10, 13, 21, 22] or a Gel’fand–Levitan integral equation [8, 21, 22].
In the reflectionless case, we can formulate the inverse problem for (2.8) as a Riemann–Hilbert problem in the complex $k$-plane. Thus, the transmission coefficients and the bound-state dependency constants are used as input to the Riemann-Hilbert problem, and the potentials are recovered from the solution to the Riemann-Hilbert problem. We refer the reader to Section 4 of [3] for the details of the analysis of the formulation of the corresponding Riemann–Hilbert problem. In this section, we present the relevant Riemann–Hilbert problem when a specific left transmission coefficient is used as input. The particular choice of the left transmission coefficient is motivated by the goal of obtaining soliton solutions to the Sawada–Kotera equation (2.1). We recall that we consider the reflectionless case for (2.8).
@@
$\Gamma(k):=\displaystyle\prod_{j=1}^{\mathbf{N}}(k+k_{j})(k+k_{j}^{\ast}).$ (5.2)
@@
We refer the reader to Section 4 of [3] for the formulation of the Riemann–Hilbert problem in the presence of reflection coefficients. Here, we present the basic steps of the formulation of the Riemann-Hilbert problem in the reflectionless case and we also describe the basic steps to solve that Riemann–Hilbert problem. For the details of the steps outlined here, we refer the reader also to Section 4 of [3]. Using the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ introduced in Section 3, we form the solutions $\Phi_{+}(k,x)$ and $\Phi_{-}(k,x)$ to (2.8) as
@@
When (5.1), (5.2), and (5.4) hold, it is known [3] that, for each fixed $x\in\mathbb{R},$ the quantity $\Phi_{+}(k,x)$ is meromorphic in $k\in\mathcal{P}^{+}$ with simple poles located at $k=k_{j}$ and $k=k_{j}^{\ast}.$ On the other hand, for each fixed $x\in\mathbb{R},$ the quantity $\Phi_{-}(k,x)$ is analytic in $k\in\mathcal{P}^{-}.$ From (5.7) we conclude that $\Phi_{+}(k,x)$ is the meromorphic continuation of $\Phi_{-}(k,x)$ from $k\in\mathcal{P}^{-}$ to $k\in\mathcal{P}^{+},$ and that $\Phi_{-}(k,x)$ is the analytic continuation of $\Phi_{+}(k,x)$ from $k\in\mathcal{P}^{+}$ to $k\in\mathcal{P}^{-}.$ Multiplying both sides of (5.7) by $e^{-kx}\,\Gamma(-k),$ we get
@@
From (5.9) and the second line of (5.5), we get
@@
We remark that (5.23) comprises a linear algebraic system of $2\mathbf{N}$ equations for the $2\mathbf{N}$ unknowns $A_{j}(x)$ with $0\leq j\leq 2\mathbf{N}-1.$ We write that linear algebraic system in the matrix form as
@@
where we have defined the $2\mathbf{N}\times 2\mathbf{N}$ matrix $\mathbf{M}(x)$ and the column vector $\mathbf{B}(x)$ with $2\mathbf{N}$ components as
@@
$m_{l}(k_{j}^{\ast})=m_{l}(k_{j})^{\ast},\qquad 1\leq j\leq\mathbf{N}.$
@@
From (5.28) we recover $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ as
@@
where the row vector in (5.29) has $1$ in the first entry and the zeros in each of the remaining $2\mathbf{N}-1$ entries. Similarly, the row vector in (5.30) has $1$ in the second entry and the zeros in all remaining entries. By using (15) on p.12 of [9], from (5.29) and (5.30) we obtain $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ as the ratio of two determinants as
@@
$A_{2\mathbf{N}-1}(x)=-\displaystyle\frac{\det[\mathbf{M}_{1}(x)]}{\det[\mathbf{M}(x)]},\qquad\mathbf{N}\geq 1,$ (5.31)
$A_{2\mathbf{N}-2}(x)=-\displaystyle\frac{\det[\mathbf{M}_{2}(x)]}{\det[\mathbf{M}(x)]},\qquad\mathbf{N}\geq 1.$ (5.32)
@@
Having determined $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ uniquely by using the input data set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ we proceed to determine the potentials $Q$ and $P$ appearing in (2.8). Using (5.2), (5.11), and (5.13) we obtain the asymptotics of $f(k,x)$ as $k\to\infty\text{\rm{ in }}\overline{\Omega_{1}}$ as
@@
$\det[\mathbf{M}_{1}(x)]=\Sigma_{\mathbf{N}}\det[\mathbf{M}(x)]+\displaystyle\frac{d}{dx}\det[\mathbf{M}(x)],$ (5.43)
@@
$A_{2\mathbf{N}-1}(x)=-\Sigma_{\mathbf{N}}-\displaystyle\frac{1}{\det[\mathbf{M}(x)]}\displaystyle\frac{d}{dx}\det[\mathbf{M}(x)].$ (5.44)
@@
As seen from (5.22) and (5.25)–(5.27), the appearance of the parameters $x$ and $t$ is restricted through the quantities $\chi_{j}$ appearing in (5.20). In fact, from (5.25)–(5.27) we observe that each of the determinants of the matrices $\mathbf{M}(x),$ $\mathbf{M}_{1}(x),$ and $\mathbf{M}_{2}(x)$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ Because of the specific appearance of $k_{j}$ and $k_{j}^{\ast}$ in (5.2) for $1\leq j\leq\mathbf{N},$ the variables $\chi_{1},\dots,\chi_{\mathbf{N}}$ appear on those three determinants in a symmetrical way. Each variable $\chi_{j}$ appears in the polynomial at most quadratically. The aforementioned symmetry allows us to determine the coefficients in those polynomials explicitly in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}}.$ For example, if we determine the coefficient of the term $\chi_{1}$ in the polynomial, then we also get the coefficient of each $\chi_{j}$ as a result of the aforementioned symmetry. By setting $\chi_{j}=0$ for $2\leq j\leq\mathbf{N},$ from (5.25) we get
@@
where the coefficients $\alpha_{0},$ $\alpha_{11},$ and $\alpha_{12}$ are explicitly expressed in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}}.$ The coefficient $\alpha_{0}$ is obtained by letting $\chi_{1}=0$ in (5.46), from which we obtain $\alpha_{0}$ as
@@
with the Vandermonde coefficient on the right-hand side of (5.49) is defined as
@@
and we recover the coefficient $\alpha_{11}$ appearing in (5.47) from (5.46) as
$\alpha_{11}=\gamma_{1}\,\mathcal{U}_{10}+\gamma_{1}^{\ast}\,\mathcal{U}_{01}.$ (5.52)
We also obtain the coefficient $\alpha_{12}$ in (5.47) as
@@
With the help of the modified Vandermonde coefficient, we get
@@
As mentioned already, the determinant of $\mathbf{M}(x)$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The coefficient of the highest-degree term $\chi_{1}^{2}\chi_{2}^{2}\cdots\chi_{\mathbf{N}}^{2}$ in $\det[\mathbf{M}(x)]$ is obtained in a manner similar to the evaluation of $\alpha_{12}$ given in (5.53), and that coefficient is given by
@@
We remark that $\mathcal{U}_{j0}$ for $2\leq j\leq\mathbf{N}$ is defined in a similar way $\mathcal{U}_{10}$ appearing in the second line of (5.54) is defined, i.e. by replacing $k_{j}$ appearing on the right-hand side of (5.49) by $zk_{j}.$ Similarly, $\mathcal{U}_{0j}$ is defined in a similar way $\mathcal{U}_{01}$ appearing in the third line of (5.54) is defined, i.e. by replacing $k_{j}^{\ast}$ appearing on the right-hand side of (5.49) by $z^{2}k_{j}^{\ast}.$ Likewise, $\mathcal{U}_{jj}$ is defined in a similar way $\mathcal{U}_{11}$ appearing in the fourth line of (5.54) is defined, i.e. by replacing $k_{j}$ and $k_{j}^{\ast}$ appearing on the right-hand side of (5.49) by $zk_{j}$ and $z^{2}k_{j}^{\ast},$ respectively.
@@ 6 Soliton solutions to the Sawada–Kotera equation
In Section 5 we have constructed complex-valued solution pairs $(Q,P)$ to the integrable system (2.6). This is done by solving the inverse scattering problem for (2.8) in the reflectionless case using the bound-state poles of the left transmission coefficient $T_{\text{\rm{l}}}(k)$ given in (5.1) and the time-evolved modified bound-state dependency constants. By using the input data set $\{k_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ with $k_{j}$ being the quantity in (5.4) and $\gamma_{j}$ being the complex-valued modified bound-state dependency constant in (5.15), we have constructed the potential pair $(Q,P)$ as in (5.39) and (5.40). Since we have used the time evolution for the dependency constants as specified in (4.5) and (5.18), the resulting time-evolved potential pair $(Q,P)$ solves the integrable coupled system (2.6).
In this section, by imposing the appropriate restriction on the solution pair $(Q,P)$ appearing in (5.39) and (5.40), we construct the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1). It turns out that relevant restriction is accomplished by specifying each of the ratios $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N}$ in the input data set $\{k_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}}$ appropriately in terms of the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in the set $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ Here, we use $r_{j}$ and $s_{j}$ to denote the real and imaginary parts of $\gamma_{j},$ respectively. Thus, we have
@@
Because of the symmetrical appearance of the quantities $k_{j}$ in (5.2), for the determination of $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N},$ it is sufficient to determine the explicit expression for $s_{1}/r_{1}$ and use the appropriate symmetries to determine the remaining ratios $s_{j}/r_{j}.$ The restrictions on the ratios $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N}$ assure that $Q$ is real valued and we have either $P\equiv 0$ or $P=Q_{x}.$ The resulting real-valued expression for $Q$ contains the $2\mathbf{N}$ real-valued parameters $\eta_{j}$ and $r_{j}$ for $1\leq j\leq\mathbf{N},$ where we recall that $\eta_{j}$ is related to $k_{j}$ as in (5.4).
@@
    $\begin{split}\left(\Pi_{\mathbf{N}}-\Sigma_{\mathbf{N}}^{2}\right)\det[\mathbf{M}(x)]+\Sigma_{\mathbf{N}}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}+&\displaystyle\frac{d^{2}\det[\mathbf{M}(x)]}{dx^{2}}+\det[\mathbf{M}_{2}(x)]\\ &=\displaystyle\frac{1}{2\,\det[\mathbf{M}(x)]}\left[\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right]^{2}.\end{split}$ (6.3)
@@
    $\begin{split}\left(\Pi_{\mathbf{N}}-\Sigma_{\mathbf{N}}^{2}\right)\det[\mathbf{M}(x)]+\Sigma_{\mathbf{N}}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}+&\displaystyle\frac{d^{2}\det[\mathbf{M}(x)]}{dx^{2}}+\det[\mathbf{M}_{2}(x)]\\ &=-\displaystyle\frac{1}{2\,\det[\mathbf{M}(x)]}\left[\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right]^{2}.\end{split}$ (6.6)
@@
    where $\alpha_{0}$ is the constant appearing in (5.47)–(5.49) and $\Delta(x)$ is a polynomial of degree $\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The use of $\alpha_{0}$ in (6.7) ensures that $\Delta(x)$ is normalized in the sense that $\Delta(x)=1$ when we choose $\chi_{j}=0$ for $1\leq j\leq\mathbf{N}.$ Comparing (5.47) with (6.7), we see that, if we set $\chi_{j}=0$ for $2\leq j\leq\mathbf{N},$ then $\Delta(x)$ must satisfy
@@
    and hence the coefficients $\alpha_{0},$ $\alpha_{11},$ and $\alpha_{12}$ appearing in (5.47) must be related to each other as
@@
    $4\,\alpha_{0}\,\alpha_{12}=\left[\alpha_{11}\right]^{2}.$ (6.10)
@@
    We divide both sides of (6.12) by $r_{1}^{2}$ and write it in the equivalent form expressed as a quadratic equation in $s_{1}/r_{1}$ as
@@
    From (5.50) and (5.55) it follows that the denominator on the right-hand side is real and nonzero, the quantity inside the square root in the numerator is positive, and $(\mathcal{U}_{10}^{2}-\mathcal{U}_{01}^{2})$ is purely imaginary. Hence, the quadratic equation (6.13) has two distinct real roots specified in (6.14). In fact, one of them is related to the constraint regarding the right-hand side of (6.3) and the other is related to the constraint regarding the right-hand side of (6.6).
@@
    Exploiting the symmetries existing in (5.47) and (5.58), with the help of (5.59), from (6.14) we obtain the ratios $s_{j}/r_{j}$ as
@@
    We already know that each of $\mathcal{U}_{00},$ $\mathcal{U}_{j0},$ $\mathcal{U}_{0j},$ $\mathcal{U}_{jj}$ is determined in terms of the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ due to (5.4). From (6.15) we observe that each ratio $s_{j}/r_{j}$ is uniquely expressed in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ in each of the cases $P\equiv 0$ and $P=Q_{x}.$ Thus, each of the quantities $(1+is_{j}/r_{j})$ is a constant uniquely expressed in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ Let us write (6.1) as

### v2

@@ 5 The inverse scattering problem in the reflectionless case
In this section we present the basic results for the inverse problem for (2.8) in the reflectionless case. In that case, the inverse problem consists of the determination of the potentials $Q$ and $P$ when the transmission coefficients $T_{\text{\rm{l}}}(k)$ and $T_{\text{\rm{r}}}(k)$ are known. In fact, as a result of (3.6), the right transmission coefficient $T_{\text{\rm{r}}}(k)$ is known when $T_{\text{\rm{l}}}(k)$ is specified. In general, the potentials $Q$ and $P$ cannot be determined uniquely when we only know the transmission coefficients. For the unique recovery of the potentials, we need some additional information. The specification of a dependency constant or a normalization constant at each bound state is a conventional way to recover the potentials uniquely. A bound-state dependency constant and a bound-state normalization constant are related [3] to each other via the residue of a transmission coefficient at the bound-state pole. The use of bound-state dependency constants is convenient when the potentials are recovered from the solution to a related Riemann-Hilbert problem. The use of bound-state normalization constants is appropriate when the potentials are recovered from the solution to a related linear integral equation, such as a Marchenko integral equation [7, 8, 10, 13, 21, 22] or a Gel’fand–Levitan integral equation [8, 21, 22].
In the reflectionless case, we formulate the inverse problem for (2.8) as a Riemann–Hilbert problem in the complex $k$-plane. Thus, the left transmission coefficient and the bound-state dependency constants can be used as input to the Riemann-Hilbert problem, and the potentials are uniquely recovered from the solution to the Riemann-Hilbert problem. We refer the reader to Section 4 of [3] for the details of the analysis of the formulation of the corresponding Riemann–Hilbert problem. In this section, we present the relevant Riemann–Hilbert problem when a specific left transmission coefficient is used as input. The particular choice of the left transmission coefficient is motivated by the goal of obtaining soliton solutions to the Sawada–Kotera equation (2.1).
@@
$\Gamma(k):=\displaystyle\prod_{j=1}^{\mathbf{N}}(k+k_{j})(k+k_{j}^{\ast}),\qquad k\in\mathbb{C}.$ (5.2)
@@
We refer the reader to Section 4 of [3] for the formulation of the Riemann–Hilbert problem in the presence of reflection coefficients. Here, we present the basic steps of the formulation of the Riemann-Hilbert problem in the reflectionless case and we also describe the basic steps to solve that Riemann–Hilbert problem. For the details of the steps outlined here, we also refer the reader to Section 4 of [3]. Using the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ introduced in Section 3, we form the solutions $\Phi_{+}(k,x)$ and $\Phi_{-}(k,x)$ to (2.8) as
@@
When (5.1), (5.2), and (5.4) hold, it is known [3] that, for each fixed $x\in\mathbb{R},$ the quantity $\Phi_{+}(k,x)$ is meromorphic in $k\in\mathcal{P}^{+}$ with simple poles located at $k=k_{j}$ and $k=k_{j}^{\ast}$ for $1\leq j\leq\mathbf{N}.$ On the other hand, for each fixed $x\in\mathbb{R},$ the quantity $\Phi_{-}(k,x)$ is analytic in $k\in\mathcal{P}^{-}.$ From (5.7) we conclude that $\Phi_{+}(k,x)$ is the meromorphic continuation of $\Phi_{-}(k,x)$ from $k\in\mathcal{P}^{-}$ to $k\in\mathcal{P}^{+},$ and that $\Phi_{-}(k,x)$ is the analytic continuation of $\Phi_{+}(k,x)$ from $k\in\mathcal{P}^{+}$ to $k\in\mathcal{P}^{-}.$ Multiplying both sides of (5.7) by $e^{-kx}\,\Gamma(-k),$ we get
@@
From (5.9) and the second line of (5.5), we obtain
@@
We remark that (5.22) and (5.23) comprise a linear algebraic system of $2\mathbf{N}$ equations for the $2\mathbf{N}$ unknowns $A_{j}(x)$ with $0\leq j\leq 2\mathbf{N}-1.$ We write that linear algebraic system in the matrix form as
@@
where we have defined the $2\mathbf{N}\times 2\mathbf{N}$ matrix $\mathbf{M}(x)$ and the column vector $\mathbf{B}(x)$ with $2\mathbf{N}$ entries as
@@
$m_{l}(k_{j}^{\ast})=m_{l}(k_{j})^{\ast},\qquad 1\leq j\leq\mathbf{N},\quad 0\leq l\leq 2\mathbf{N}.$
@@
From (5.28) we recover $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x),$ respectively, as
@@
where the row vector in (5.29) has $1$ in its first entry and the zeros in each of the remaining $2\mathbf{N}-1$ entries. Similarly, the row vector in (5.30) has $1$ in its second entry and the zeros in all the remaining entries. By using (15) on p.12 of [9], from (5.29) and (5.30) we obtain $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ as the ratio of two determinants as
@@
$A_{2\mathbf{N}-1}(x)=-\displaystyle\frac{\det[\mathbf{M}_{1}(x)]}{\det[\mathbf{M}(x)]},$ (5.31)
$A_{2\mathbf{N}-2}(x)=-\displaystyle\frac{\det[\mathbf{M}_{2}(x)]}{\det[\mathbf{M}(x)]}.$ (5.32)
@@
Having determined $A_{2\mathbf{N}-1}(x)$ and $A_{2\mathbf{N}-2}(x)$ uniquely by using the input data set $\{\eta_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ we proceed to determine the potentials $Q$ and $P$ appearing in (2.8). Using (5.2), (5.11), and (5.13), we obtain the asymptotics of $f(k,x)$ as $k\to\infty\text{\rm{ in }}\overline{\Omega_{1}}$ as
@@
$\det[\mathbf{M}_{1}(x)]=\Sigma_{\mathbf{N}}\det[\mathbf{M}(x)]+\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx},$ (5.43)
@@
$A_{2\mathbf{N}-1}(x)=-\Sigma_{\mathbf{N}}-\displaystyle\frac{1}{\det[\mathbf{M}(x)]}\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}.$ (5.44)
@@
As seen from (5.22) and (5.25)–(5.27), the appearance of the parameters $x$ and $t$ is restricted through the quantities $\chi_{j}$ appearing in (5.20). In fact, from (5.25)–(5.27) we observe that each of the determinants of the matrices $\mathbf{M}(x),$ $\mathbf{M}_{1}(x),$ and $\mathbf{M}_{2}(x)$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ Because of the specific appearance of $k_{j}$ and $k_{j}^{\ast}$ in (5.2) for $1\leq j\leq\mathbf{N},$ the variables $\chi_{1},\dots,\chi_{\mathbf{N}}$ appear on those three determinants in a symmetrical way. Each variable $\chi_{j}$ appears in the polynomial at most quadratically. The aforementioned symmetry allows us to determine the coefficients in those polynomials explicitly in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}}.$ For example, if we determine the coefficient of the term $\chi_{1}$ in the polynomial, then we also get the coefficient of each $\chi_{j}$ by exploiting the aforementioned symmetry. By setting $\chi_{j}=0$ for $2\leq j\leq\mathbf{N},$ from (5.25) we get
@@
where the coefficients $\alpha_{0},$ $\alpha_{11},$ and $\alpha_{12}$ are explicitly expressed in terms of the parameters in the set $\{k_{j},k_{j}^{\ast},\gamma_{j},\gamma_{j}^{\ast}\}_{j=1}^{\mathbf{N}}.$ The coefficient $\alpha_{0}$ is obtained by letting $\chi_{1}=0$ in (5.46), from which we have
@@
where the Vandermonde coefficient on the right-hand side of (5.49) is defined as
@@
we recover the coefficient $\alpha_{11}$ appearing in (5.47) from (5.46) as
$\alpha_{11}=\gamma_{1}\,\mathcal{U}_{10}+\gamma_{1}^{\ast}\,\mathcal{U}_{01},$ (5.52)
and we also obtain the coefficient $\alpha_{12}$ in (5.47) as
@@
With the help of the modified Vandermonde coefficient, we determine the coefficients $\beta_{0},$ $\beta_{11},$ $\beta_{12}$ appearing in (5.55) as
@@
As mentioned already, the determinant of $\mathbf{M}(x)$ is a polynomial of degree $2\mathbf{N}$ in the $\mathbf{N}$ real-valued variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The coefficient of the highest-degree term $\chi_{1}^{2}\chi_{2}^{2}\cdots\chi_{\mathbf{N}}^{2}$ in $\det[\mathbf{M}(x)]$ is obtained in a manner similar to the evaluation of $\alpha_{12}$ given in (5.53), and that coefficient is equal to the quantity given by
@@
We remark that $\mathcal{U}_{j0}$ for $2\leq j\leq\mathbf{N}$ is defined in a similar way $\mathcal{U}_{10}$ appearing in the second line of (5.54) is defined, i.e. by replacing $k_{j}$ appearing on the right-hand side of (5.49) by $zk_{j}.$ Similarly, $\mathcal{U}_{0j}$ is defined in an analogous manner $\mathcal{U}_{01}$ appearing in the third line of (5.54) is defined, i.e. by replacing $k_{j}^{\ast}$ appearing on the right-hand side of (5.49) by $z^{2}k_{j}^{\ast}.$ Likewise, $\mathcal{U}_{jj}$ is defined in a similar way $\mathcal{U}_{11}$ appearing in the fourth line of (5.54) is defined, i.e. by replacing $k_{j}$ and $k_{j}^{\ast}$ appearing on the right-hand side of (5.49) by $zk_{j}$ and $z^{2}k_{j}^{\ast},$ respectively.
@@ 6 Soliton solutions to the Sawada–Kotera equation
In Section 5 we have constructed complex-valued solution pairs $(Q,P)$ to the integrable system (2.6). This is done by solving the inverse scattering problem for (2.8) in the reflectionless case by using as input the bound-state poles of the left transmission coefficient $T_{\text{\rm{l}}}(k)$ given in (5.1) and the time-evolved modified bound-state dependency constants. By using the input data set $\{k_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}},$ with $k_{j}$ being the quantity in (5.4) and $\gamma_{j}$ being the complex-valued modified bound-state dependency constant in (5.15), we have constructed the potential pair $(Q,P)$ as in (5.39) and (5.40). Since we have used the time evolution for the dependency constants as specified in (4.5) and (5.18), the resulting time-evolved potential pair $(Q,P)$ constitutes a solution to the integrable coupled system (2.6).
In this section, by imposing the appropriate restriction on the solution pair $(Q,P)$ appearing in (5.39) and (5.40), we construct the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1). It turns out [3] that relevant restriction is accomplished by specifying each of the ratios $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N}$ in the input data set $\{k_{j},\gamma_{j}\}_{j=1}^{\mathbf{N}}$ appropriately in terms of the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in the set $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ Here, we use $r_{j}$ and $s_{j}$ to denote the real and imaginary parts of $\gamma_{j},$ respectively. Thus, we have
@@
Because of the symmetrical appearance of the quantities $k_{j}$ in (5.2), for the determination of $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N},$ it is sufficient to determine the explicit expression only for $s_{1}/r_{1}$ and use the appropriate symmetries to determine the remaining ratios $s_{j}/r_{j}.$ The restrictions on the ratios $s_{j}/r_{j}$ for $1\leq j\leq\mathbf{N}$ assure that $Q$ is real valued and we have either $P\equiv 0$ or $P=Q_{x}.$ The resulting real-valued expression for $Q$ contains the $2\mathbf{N}$ real-valued parameters $\eta_{j}$ and $r_{j}$ for $1\leq j\leq\mathbf{N},$ where we recall that $\eta_{j}$ is related to $k_{j}$ as in (5.4).
@@
    $\begin{split}\left(\Pi_{\mathbf{N}}-\Sigma_{\mathbf{N}}^{2}\right)\det[\mathbf{M}(x)]+\Sigma_{\mathbf{N}}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}+&\displaystyle\frac{d^{2}\det[\mathbf{M}(x)]}{dx^{2}}+\det[\mathbf{M}_{2}(x)]\\ &=\displaystyle\frac{1}{2\,\det[\mathbf{M}(x)]}\left(\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right)^{2}.\end{split}$ (6.3)
@@
    $\begin{split}\left(\Pi_{\mathbf{N}}-\Sigma_{\mathbf{N}}^{2}\right)\det[\mathbf{M}(x)]+\Sigma_{\mathbf{N}}\,\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}+&\displaystyle\frac{d^{2}\det[\mathbf{M}(x)]}{dx^{2}}+\det[\mathbf{M}_{2}(x)]\\ &=-\displaystyle\frac{1}{2\,\det[\mathbf{M}(x)]}\left(\displaystyle\frac{d\det[\mathbf{M}(x)]}{dx}\right)^{2}.\end{split}$ (6.6)
@@
    where $\alpha_{0}$ is the constant appearing in (5.47)–(5.49) and $\Delta(x)$ is a polynomial of degree $\mathbf{N}$ in $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The use of $\alpha_{0}$ in (6.7) ensures that $\Delta(x)$ is normalized in the sense that $\Delta(x)=1$ when we let $\chi_{j}=0$ for $1\leq j\leq\mathbf{N}.$ Comparing (5.47) with (6.7), we see that, if we set $\chi_{j}=0$ for $2\leq j\leq\mathbf{N},$ then $\Delta(x)$ must satisfy
@@
    and hence the coefficients $\alpha_{0},$ $\alpha_{11},$ $\alpha_{12}$ appearing in (5.47) must be related to each other as
@@
    $4\,\alpha_{0}\,\alpha_{12}=\left(\alpha_{11}\right)^{2}.$ (6.10)
@@
    We divide both sides of (6.12) by $r_{1}^{2}$ and write the resulting equality in the equivalent form expressed as a quadratic equation in $s_{1}/r_{1}$ as
@@
    From (5.50) and (5.55) it follows that the denominator on the right-hand side of (6.14) is real and nonzero, the quantity inside the square root in the numerator is positive, and $(\mathcal{U}_{10}^{2}-\mathcal{U}_{01}^{2})$ is purely imaginary. Hence, the quadratic equation (6.13) has two distinct real roots specified in (6.14). In fact, one of them is related to the constraint regarding the right-hand side of (6.3) and the other is related to the constraint regarding the right-hand side of (6.6).
@@
    By exploiting the symmetries existing in (5.47) and (5.58), with the help of (5.59), from (6.14) we obtain the ratios $s_{j}/r_{j}$ as
@@
    We already know that each of $\mathcal{U}_{00},$ $\mathcal{U}_{j0},$ $\mathcal{U}_{0j},$ $\mathcal{U}_{jj}$ is determined in terms of the elements in the set $\{k_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in the set $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ due to (5.4). From (6.15) we observe that each ratio $s_{j}/r_{j}$ is uniquely expressed in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}$ in each of the cases $P\equiv 0$ and $P=Q_{x}.$ Thus, each of the quantities $(1+is_{j}/r_{j})$ is a constant uniquely expressed in terms of the elements in the set $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ Let us write (6.1) as